OpenStudy (anonymous):

Witch is greater 4/5 or 2/3

OpenStudy (anonymous):

4/5 (80%) > 2/3 (66%) If 1/5 is .2, then .2 * 4 = .8 If 1/3 = .33... then .33... * 2 = .66...

OpenStudy (anonymous):

Not sure if it's wrong to say here, but the correct spelling is which.

sammixboo (sammixboo):

\(\bf \huge \color{magenta}{1^{st}~Way~to~Solve~This~Problem}\) Convert both of the fractions into decimals. 4/5 = ? 2/3 = ? Whichever decimal has the greater value is your answer! \(\bf \huge \color{magenta}{2^{nd}~Way~to~Solve~This~Problem}\) You can draw out the fractions as pictures, and compare them to see which has a greater amount. Let's draw rectangles.